Sec^2 x - 2tan x = 0
Answer by stanbon(75887)

Sec^2 x - 2tan x = 0
----
tan^2(x)+1 - 2tan(x) = 0
----
tan^2 - 2tan + 1 = 0
-------
(tan-1) = 0
-------
tan(x) = 1
x = pi/4 or (5/4)pi
---------------------
Cheers,
Stan H.
-------------